(1) Alex Company manufactures boats in bottles (to sell to rich people who don’t have the time to make them themselves)....

(1)

Alex Company manufactures boats in bottles (to sell to rich people who don’t have the time to make them themselves). For their standard costing system they allow 4 hours of direct labor per bottle at a wage of $12 per hour. The last batch only took 4.5 hours per bottle. They made 12 bottles. However, they ended up paying employees $12 per hour. What is the standard cost for labor per bottle? (in dollars per bottle)

(2)

Garland Company uses a standard cost system. The standard for each finished unit of product allows for 2 pounds of plastic at $0.71 per pound. During December, Garland bought 4000 pounds of plastic at $0.76 per pound, and used 4100 pounds in the production of 1200 finished units of product. What is the direct materials purchase price variance for the month of December?

Please indicate an unfavorable or favorable variance with "u" or "f", respectively.

(3)

NewTone, Inc makes MP3 players and uses standard costing. They recently used 20,000 labor hours to produce 8,000 units. They originally budgeted 20,000 hours to produce 10,000 units with payroll of $333,000. The company’s actual payroll cost amounted to $300,000. What was the direct labor efficiency variance?

Please round to the nearest dollar at the end of your calculations. Also, type "u" for unfavorable or "f" for favorable to indicate the direction of the variance.
Answer #1

1) Standard time to make 1 bottle is 4 Hrs

Labour wages are paid @ $12 per hour

Hence, Standard labour cost per bottle = 12×4 = $48 per bottle.

2) Direct Material Purchase price variance = AQ( SP -AP)

where,

AQ = Actual quantity of the material purchased= 4000 pounds

SP= Standard Price per unit of material=$ 0.71 per pound

AP = Actual price per unit of material= $ 0.76 per pound

Hence,

variance = 4000(0.71-0.76)

= 200(u)

3) Direct Labour Efficiency Variance= SR(SH-AH paid)

where, SR= Standard rate per labour hour= $ 333000/20000 hours

= $ 16.65 per labour hour

SH = Standard hours required for actual output

• For 10000 units 20000 hours are required

so for 8000 units ,standard hours required are 20000/10000× 8000

= 16000 hours

AH paid= Actual Labour hours paid= 20000 hours

Hence variance = 16.65( 16000-20000)

= 16.65×(-4000)

= 66600(u)